The density of gold is 19.3 g/cm3. A gold dealer advertized a bar of pure gold for sale. The Gold bar has a mass of 3300.00 g and measured 2.00 cm x 15.0 cm x 6.00 cm. Was the bar pure gold
1
Expert's answer
2013-01-15T08:22:58-0500
Thedensity of gold is 19.3 g/cm3. A gold dealer advertized a bar of pure gold for
sale. The Gold bar has a mass of 3300.00 g and measured 2.00 cm x 15.0 cm x
6.00 cm. Was the bar pure gold
V(Bar) = 2*15*6 = 180 cm3
m(gold) = 180cm3 * 19,3g/cm3 = 3474 g
pure gold = 3300/3474 =* 100% = 95 %